Project A

Project B

Initial cash outlay

$180,000

$160,000

Annual depreciation

$25,000

$20,000

Annual net cash inflow after tax

$50,000

$40,000

Expected salvage value

$0

$0

Projects A and B are to be evaluated using the payback period and the unadjusted rate of return. State which project should be accepted under each method.
Answer #1

Payback period = Initial cash outlay / Annual net cash inflow after tax

Project A = 180000 / 50000 = 3.6 years

Project B = 160000 / 40000 = 4 years

unadjusted rate of return = net income / Initial cash outlay

Project A = (50000-25000) / 180000 = 13.89%

Project B = (40000-20000)/160000 = 12.55%

Project A should be accepted under each method.